Graph the given curve and find the area of a bounded region.


y= 4x - x^2

1
Expert's answer
2022-03-14T17:35:34-0400

Solution


The required area is found by using the formula


"A = \int\limits_a^b {f\left( x \right)} \,\,dx"


Here given that


"f(x)= 4x - x^2"


When plotted as shown below, we can see, the bounded area, in this case, is under the curve "f(x)= 4x - x^2", the "x-axis" and the two limits "x=0" and "x=4"


Therefore, the required area is calculated as


"A = \int\limits_0^4 {(4x-x^2)} \,\,dx"


"\begin{array}{l}
A = \int\limits_0^4 {\left( {4x - {x^2}} \right)} \,\,dx\\
A = \left[ {\frac{{4{x^2}}}{2} - \frac{{{x^3}}}{3}} \right]_0^4\\
A = \left[ {\left( {2{{\left( 4 \right)}^2} - \frac{{{{\left( 4 \right)}^3}}}{3}} \right) - \left( {2{{\left( 0 \right)}^2} - \frac{{{{\left( 0 \right)}^3}}}{3}} \right)} \right]\\
A = \left[ {32 - \frac{{64}}{3} - 0 + 0} \right]\\
A = \frac{{32}}{3}
\end{array}"


Hence required area is "\frac{32}{3}" square units